# arrows.meta, chains, positioning

### v1

@@ VII Bayesian Optimisation Results
[ xlabel=Episode, ylabel=Rewards, legend pos=north west, legend style=font=, legend cell align=left, ymajorgrids=true, grid style=dashed, ymin=-15000, ymax=60000, scaled y ticks = false, y tick label style=/pgf/number format/sci, width=height=6cm]
[…]
@@
[ xlabel=Episode, ylabel=Rewards, legend pos=north west, legend style=font=, legend cell align=left, ymajorgrids=true, grid style=dashed, ymin=-15000, ymax=60000, scaled y ticks = false, y tick label style=/pgf/number format/sci, width=height=6cm]
[color=red,mark=x] coordinates (1, -944) (2, -667) (3, 919) (4, 4009) (5, 7860) (6, 14303) (7, 21405) (8, 28799) (9, 38095) (10, 48823) ;
[color=blue,mark=x] coordinates (1, -588) (2, 2080) (3, 3122) (4, 6573) (5, 9759) (6, 16085) (7, 25136) (8, 34720) (9, 46093) (10, 57388) ;
[color=green,mark=x] coordinates (1, 909) (2, -258) (3, 1398) (4, 5207) (5, 10434) (6, 17728) (7, 25584) (8, 34898) (9, 32468) (10, 42637) ;
[color=orange,mark=x] coordinates (1, -84) (2, 272) (3, 2355) (4, 5792) (5, 11162) (6, 17760) (7, 23002) (8, 31702) (9, 40771) (10, 52126) ;
@@ VIII Conclusion
In conclusion, this study proposed a novel GraphRL framework based on a Temporal Graphical Convolutional Network (T-GCN) to forecast the future states of a dynamic RL environment and overcome the limitations of deep learning models in time-series prediction tasks. The proposed framework was fine-tuned with the Bayesian optimisation technique and was applied to three datasets: WESAD, LAM Traffic Forecasting, and US Weather Forecasting. The results showed that the proposed GraphRL agent outperformed other traditional RL agents in all three predictive and monitoring applications. The agent’s performance on the WESAD dataset was relatively consistent, while on the LAM Traffic Forecasting and US Weather Forecasting datasets, the agent’s performance was relatively inconsistent. The inconsistency of the scores was due to the exploration rate. The Bayesian optimisation process was able to fine-tune the $L_{reg}$, $\lambda$ and minimise the loss function to improve the performance of the GraphRL agent. The proposed GraphRL framework can be applied to a wide range of applications and can be further enhanced by incorporating other forms of information such as node and edge attributes for more accurate predictions.
Ethical Statement
There are no ethical issues.

### v2

@@ VIII Conclusion
The GraphRL framework, introduced in this study, embodies an innovative amalgamation of T-GCN and RL. It is specifically engineered to augment the prediction of future states in dynamic environments. Rigorous evaluations, utilizing an array of datasets such as WESAD, LA Traffic Forecasting, and US Weather Forecasting, have substantiated the framework’s enhanced performance compared to conventional RL agents. Nonetheless, it is imperative to acknowledge that the efficacy of GraphRL is significantly contingent upon the caliber of the input data and necessitates substantial computational resources. The framework’s reliance on data of high quality and structure constitutes a considerable limitation, with its accuracy and effectiveness being closely bound to the data’s integrity. Additionally, the computational requisites, predominantly due to the T-GCN model integration, present challenges in scalability and broader applicability.
Future enhancements of the GraphRL framework will be directed towards surmounting these constraints and broadening its functional scope. Prospective developments entail the incorporation of spatial data processing, aimed at bolstering the framework’s analytical prowess, particularly in processing data with spatial or geographical dimensions. Investigating a spectrum of graph-based models could yield insights for enhancing both the efficiency and efficacy of the framework. Furthermore, the exploration of real-time adaptive learning strategies presents a promising avenue for subsequent research. Such advancements are anticipated to enable the framework to dynamically adapt to evolving data patterns and environmental shifts. In summation, the GraphRL framework signifies a substantial advancement in the domain of time-series prediction and monitoring. Its adeptness in managing complex temporal data surpasses traditional RL methodologies, heralding innovative applications in sectors such as healthcare, traffic management, and environmental forecasting. As the framework undergoes continued refinement and evolution, it is positioned to emerge as an instrumental component in the progression of predictive analytics and intelligent monitoring systems, with extensive applicability across diverse fields.